OpenStudy (anonymous):

integrate t.e^(-0.1t)

OpenStudy (dumbcow):

looks like integration by parts problem \[\int\limits udv = uv - \int\limits vdu\] u = t dv = e^(-.1t) du = dt v = -10e^(-.1t)